Identified - This is a repost after previous incident was closed.
Update - Maintenance work will continue starting Wednesday, Jan 7th, on a cable that serves Walker, Everson and Hart Halls.
On Nov 3, 12:40 PST it was reported a problem with telephone service interruption. the NOC personnel identified a problem moisture in the cable splice cases that affected Wickson, Hunt, Hart, Freeborn and Viehmeyer buildings.
Jan 15, 12:12 PST
Identified - Work on a cable splice will begin Tuesday, Jan 6th. Some users may experience intermittent static or loss of
service. The following buildings may be affected:
Veg Crops Bulb Storage House
Plant Reproductive Biology Facility
Bowley Plant Science Teaching Facility
Hutchison Child Development Center
1050 Extension Center Drive
Recreation Pool Lodge
Jan 5, 15:34 PST
Purpose: Upgrading code on the wireless controller for these areas to the vendor-recommended version.
Impact: Loss of wireless network connectivity as the controller is rebooted.
Locations Affected: The following buildings will be affected:
CUARTO (OXFORD) COMMONS
PRIMERO GROVE COMMUNITY
PRIMERO GROVE LAUREL
PRIMERO GROVE MAGNOLIA
PRIMERO GROVE MANZANITA
PRIMERO GROVE SPRUCE
REGAN CENT C
IT Express Service Desk
Jan 22, 12:43 PST
Maintenance on the DaFIS-R1 Server has been completed.
Jan 27, 16:12 PST
The hard drive is expected to be replaced Tuesday Jan 27.
Jan 26, 07:08 PST
Starting around 16:45 Friday, a hard drive triggered a fault notice on the DaFIS-R1 Server. A reboot was attempted but the DaFIS-R1 Server will not come back up on its own. Maintenance on the drive is expected to continue throughout the weekend (48-72 hours).
Jan 23, 21:26 PST
The scheduled maintenance has been completed.
Jan 26, 07:00 PST
In progress -
Scheduled maintenance is currently in progress. We will provide updates as necessary.
Jan 26, 06:00 PST
Purpose: Applying vendor recommended system configuration changes to the primary Campus Procera system, and moving traffic back onto the primary path to test and verify that traffic latency issue is resolved.
Impact: Possibility of increased latency for traffic to/from the Internet and I2 if configuration changes do not resolve the issue as expected. If latency is still present, traffic will be moved back to secondary path.
Affected: All traffic to/from the Internet and I2.
IT Express Service Desk
Jan 21, 11:02 PST